OpenStudy (anonymous):

Use the distributive property to simplify the expression below: The quantity of negative 3 times x plus 9 all divided by negative 3 -x - 3 x + 3 x + 9 x - 3

OpenStudy (anonymous):

(3x+9)/-3

OpenStudy (anonymous):

Thank you soo much =)

OpenStudy (anonymous):

wait lol yes thats what the problem is

OpenStudy (anonymous):

no problem, I was already doing the next step, but glad you got it :)

OpenStudy (anonymous):

so we continue?

OpenStudy (anonymous):

now that we have the equation and see how it looks we notice that we can factor the numerator.

OpenStudy (anonymous):

lol yes

OpenStudy (anonymous):

lol :p... could you factor the numerator?

OpenStudy (anonymous):

ok so i forgot ho to do this so is it 12x/3 which is 6?

OpenStudy (anonymous):

um, where did you get the 12 from?

OpenStudy (anonymous):

i added 3x+9 lol

OpenStudy (anonymous):

oh lol yea.. you are off :p

OpenStudy (anonymous):

smh ohh thats sucks lol

OpenStudy (anonymous):

Firstly remember that when adding things in an equation you can only add things with common variables

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

when factoring we are looking to simplify what we have in front of us. a common coefficient that both numbers have in numerator is 3 so 3(x+9)=3x+9

OpenStudy (anonymous):

wait i messed up there

OpenStudy (anonymous):

3x+9=3(x+3)

OpenStudy (anonymous):

if we distribute the 3 we get back to 3x+9

OpenStudy (anonymous):

ok i get it

OpenStudy (anonymous):

cool, you know what to do next?

OpenStudy (anonymous):

yes thank you sooo much =)

OpenStudy (anonymous):

awesome! :D and you are welcomed